# A Compact Analytical Solution of the Dicke Superradiance Master Equation via Residue Calculus

### v6

@@ Abstract
This operator acts within the fully symmetric subspace of the total Hilbert space, spanned by Dicke states |$\ket{m}$ with excitation numbers $m=0,1,\dots,N$. The ladder structure is defined by
$S\ket{m}=\sqrt{h_{m}}\ket{m-1},\quad S^{\dagger}\ket{m}=\sqrt{h_{m+1}}\ket{m+1},$
@@
where $\bar{m}=N+1-m$ and the notation $[\cdot]_{z=h_{j}}$ indicates that the expression is evaluated at $z=h_{j}$. The Heaviside functions $\theta_{m>N/2}$ and $\theta_{m\leq N/2}$ are defined to be 1 if the condition is satisfied and 0 otherwise. These functions distinguish between the non-degenerate region above the equator (where $h_{m}\neq h_{\bar{m}}$) and the degenerate region below it (where $h_{m}=h_{\bar{m}}$).
@@
Figure 1: Left: Contributions to the solution in Eq. (3) above the equator adding only simple poles marked in blue. Right: Separation of simple pole contributions in blue and double poles contributions ($h_{j}=h_{\bar{j}}$) marked in red, for even $N$.
@@
To begin, we express the time-evolved density operator using a power series expansion in the Lindbladian superoperator $\mathcal{L}$. Applied to the initial condition $\rho(0)=\ket{N}\bra{N}$ |, the evolution reads:
@@
where $h_{m}=m(N+1-m)$ are the familiar Dicke ladder factors. This identity reflects the fact that the system may either remain in the current Dicke state (with weight $-h_{m}$) or decay to the next lower excitation level (with weight $h_{m}$).
Since the bottom state |$\ket{0}$ has no further decay, we find [|$\mathcal{L}[\ket{0}\bra{0}]=0$. Using this recursive structure, we can compute higher-order corrections $\rho_{m}^{(j)}$ by applying the following recurrence relation:
@@
$\rho(0)=\ket{m_{0}}\bra{m_{0}},\quad\text{with }1\leq m_{0}\leq N.$
This scenario corresponds to an ensemble that starts with only $m_{0}$ excitations rather than all $N$. Since the system evolves strictly downward in excitation number due to collective decay, the accessible Hilbert space is now truncated: only Dicke states |$\ket{m}$ with $0\leq m\leq m_{0}$ will be populated during the evolution.
@@
$\rho(0)=\sum_{k=0}^{N}\rho_{k}(0)\ket{k}\bra{k}.$ | .
Using linear superposition, the full solution is simply the weighted sum of individual evolutions from each pure component |$\ket{k}$:
@@
Acknowledgments - We acknowledge financial support from the Max Planck Society and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 429529648 – TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”). R.H. acknowledges funding by the Austrian Science Fund (FWF) 10.55776/W1259.

### v7

@@ Abstract
[…]
@@
where $\bar{m}=N+1-m$ and the notation $\cdot]_{z=h_{j}}$ indicates that the expression is evaluated at $z=h_{j}$. The Heaviside functions $\theta_{m>N/2}$ and $\theta_{m\leq N/2}$ are defined to be 1 if the condition is satisfied and 0 otherwise. These functions distinguish between the non-degenerate region above the equator (where $h_{m}\neq h_{\bar{m}}$) and the degenerate region below it (where $h_{m}=h_{\bar{m}}$).
@@
Figure 1: Left: Contributions to the solution in Eq. (3) above the equator adding only simple poles marked in blue. Right: Separation of simple pole contributions in blue and double poles contributions ($h_{j}=h_{\bar{j}}$) marked in red, for even NN.
@@
To begin, we express the time-evolved density operator using a power series expansion in the Lindbladian superoperator $\mathcal{L}$. Applied to the initial condition $|\rho(0)=\ket{N}\bra{N}$, the evolution reads:
@@
where $h_{m}=m(N+1-m)$ are the familiar Dicke ladder factors. This identity reflects the fact that the system may either remain in the current Dicke state (with weight $h_{m}$) or decay to the next lower excitation level (with weight $h_{m}$).
[…]
@@
[…]
@@
$|.\rho(0)=\sum_{k=0}^{N}\rho_{k}(0)\ket{k}\bra{k}$.
[…]
@@
Acknowledgments - We acknowledge financial support from the Max Planck Society and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 429529648 – TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”). This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/COE1].